# Uniform convergence of the smooth calibration error and its relationship with functional gradient

### v2

@@ 4.1 Case study I: Gradient Boosting Tree
Under the same assumptions as in Theorem 3, there exist universal constants $\{C_{i}\}$s such that with probability at least $1-\delta$ over the draw of $S_{\mathrm{tr}}$, we have:
$\min_{t\in\{1,\dots,T\}}\mathrm{smCE}(g^{(t)},\mathcal{D})\leq\frac{C_{5}B^{2}}{T+2}+\frac{C_{4}}{\sqrt{n}}+C_{6}\sqrt{\frac{2^{m}((\log 2n)^{5}+\log d)}{n}}+\sqrt{\frac{2\log(3/\delta)}{n}}.$ (19)
To establish this result, we use the inequality $\mathrm{smCE}(f,S_{\mathrm{tr}})\leq\mathrm{smCE}^{\sigma}(g,S_{\mathrm{tr}})$, which serves as the empirical analogue of the corresponding bound at the population smooth CE. This implies that Theorem 3 also provides a guarantee for the training smooth CE. We obtain the desired bound by combining this with the uniform convergence bound in Theorem 2.
We use Theorem 2, which is based on the Rademacher complexity, as it allows tractable estimation for the Rademacher complexity of the convex hull of $\{\psi_{\theta}\}$. Using Corollary 2, we show that with suitable choices of $T$ and $n$, the smooth CE can be controlled to within a desired precision level $\epsilon$.
@@
Under the same assumptions as in Corollary 2, for any $\epsilon>0$, if $T=\Omega(\epsilon^{-1}),n=\tilde{\Omega}(\epsilon^{-2})$, then, with probability at least $1-\delta$, $\min_{t\in\{1,\dots,T\}}\mathrm{smCE}(g^{(t)},\mathcal{D})\leq\epsilon$.
Here, $\tilde{\Omega}(\cdot)$ hides logarithmic factors in the big-$\Omega$ notation. Corollary 3 establishes that $\epsilon$-smooth CE can be achieved using $\mathcal{O}(1/\epsilon)$ iterations of GBT.
We remark that the standard margin theory provides a generalization bound on the test accuracy of GBTs without requiring the depth condition $m\geq d$. This result is well-established in the literature [Zhang_2023], and for completeness, we include a formal statement in Appendix D. In conjunction with Corollary 3, this yields, to the best of our knowledge, the first theoretical analysis of GBTs that accounts for both test accuracy and smooth CE. We further support our analysis with numerical experiments in Appendix H, investigating how the number of iterations and training sample size influence the smooth CE and prediction accuracy.
@@ 4.2 Case study II: Kernel Boosting
where $\mathcal{T}_{k}\nabla_{g}L_{n}(g)\coloneqq\frac{1}{n}\sum_{i=1}^{n}k(X_{i},\cdot)\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X_{i})),Y_{i}))$ is the empirical kernel operator. In kernel boosting, we set $h_{t}=-\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})$ in Eq. (15) and use the update rule:
@@
Here, $12Y_{i}-$ 1 converts the binary label $Y_{i}\in\{0,1\}$ into ±$\pm 1$. This condition guarantees the existence of a separating function in $\mathcal{H}$ with the margin $\gamma$ for the training dataset, and is a standard assumption in the theoretical analyses of classification [wei2018margin] and boosting [telgarsky2013margins]. Under this assumption, the $L_{1}$ norm of the functional gradient is bounded as follows:
@@
All proofs for this section are provided in Appendix E. From Eq. (14), it follows that $\mathrm{smCE}^{\sigma}(g,S_{\mathrm{tr}})\leq\|\nabla_{g}\ell_{\mathrm{ent}}(g(X),Y)\|_{L_{1}(S_{n})}$. Using this, we obtain the following result.
@@
Suppose Assumption 1, $\sup_{x,x^{\prime}\in\mathcal{X}}k(x,x^{\prime})\leq\Lambda$, and $\|g^{(0)}\|_{\mathcal{H}}\leq\Lambda^{\prime}$ hold. When using the constant stepsize $w_{t}=w(<4/\Lambda)$, the average of function $\bar{g}^{(T)}=\frac{1}{T}\sum_{t=0}^{T-1}g^{(t)}$ satisifies
@@
$\mathrm{smCE}(\bar{g}^{(T)},\mathcal{D})\leq\frac{1}{\gamma}\sqrt{\frac{L_{n}(g^{(0)})}{wT}}\!+\frac{C_{7}}{\sqrt{n}}\!+C_{8}(\Lambda^{\prime}+\sqrt{2wTL(g^{(0)})})\sqrt{\frac{\Lambda}{n}}(\log n)^{2}\!+\sqrt{\frac{2\log\frac{3}{\delta}}{n}}.$ (24)
We also note that bounding the norm of the functional gradient leads to a generalization guarantee for the test accuracy [nitanda18a]. For completeness, we provide a formal upper bound on the misclassification rate ($P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]$) in Appendix E.
Although covering number bounds require assumptions on kernel eigenvalue decay, here we use the Rademacher complexity bound from Theorem 2, which yields a simpler (albeit slightly looser) result.
We observe that increasing the number of steps $T$ reduces the training smooth CE, while simultaneously increasing the capacity of the function class, as the Rademacher complexity grows with the norm of $\mathcal{H}$ at the rate $O(\sqrt{wT})$. This implies a trade-off between minimizing the training smooth CE and controlling model complexity. We show that appropriate choices of hyperparameters ensure desired levels of precision $\epsilon$ for both the smooth CE and the misclassification rate.
@@
then, with probability at least $1-\delta$, the average of function $\bar{g}^{(T)}$ satisfies $\mathrm{smCE}(\bar{g}^{(T)},\mathcal{D})\leq\epsilon$ and $P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq\epsilon$.
@@
Remark 2.
Unlike kernel boosting, Theorem 3 for GBTs does not require a margin assumption. This difference stems from the method of the analysis: in kernel boosting, we upper bound the $L_{1}$ norm of the functional gradient (Eq. (22)), whereas in Theorem 3 for GBT, we directly bound the inner product in Eq. (14) using structural properties of trees, without relying on the functional gradient norm.
@@ 4.3 Case study III: Two-layer neural network
The next example is the two-layer neural network (NN). Whereas NNs are typically trained via gradient descent on their parameters, a recent work [nitanda2019gradient] has shown that, under certain hyperparameter settings—such as output scaling and a sufficiently large number of neurons—two-layer NNs behave similarly to kernel boosting algorithms associated with the neural tangent kernel (NTK).
We follow the settings of the analysis by nitanda2019gradient: Let the logit functon $g_{\theta}:\mathcal{X}\rightarrow\mathbb{R}$ be defined by a two-layer NN as $g_{\theta}(x)=\frac{1}{m^{\beta}}\sum_{r=1}^{m}a_{r}\phi(\theta_{r}\cdot x)$, where $m$ is the number of hidden units, $\beta$ is a scaling exponent, and $\phi:\mathbb{R}\to\mathbb{R}$ is a smooth activation function (e.g., sigmoid, tanh). The weights $\{a_{r}\}_{r=1}^{m}\in\{-1,1\}^{m}$ are fixed, while the parameters $\theta=\{\theta_{r}\}_{r=1}^{m}$, with $\theta_{r}\in\mathbb{R}^{d}$, are updated via gradient descent: $\theta^{(t+1)}=\theta^{(t)}-w\nabla_{\theta}L_{n}(g_{\theta^{(t)}})$, where $w>0$ is a constant stepsize.
Here, we present the informal statement of Theorem 2 by nitanda2019gradient, which provides the guarantee of the functional gradient: Assume that 1) the activation function is smooth, 2) the initial parameters are sampled from a sub-Gaussian distribution, 3) the transformed data by NTK is separable with the margin $\gamma$, 3) the stepsize is sufficiently small, and 4) the number of hidden units is sufficiently large $m$. Then, for $T\leq\frac{m\gamma^{2}K_{3}}{w}$, there exist the constants $K_{1}$, $K_{2}$, and $K_{3}$ such that, for all $\beta\in(0,1]$, with probability at least $1-\delta$ over the random initialization, the following bound holds:
@@
See Appendix F for the formal statement. Then, similar to Theorem 4, we have
@@
Combining the covering number estimate by nitanda2019gradient with our Theorem 2, we obtain the upper bounds on the population smooth CE and misclassification rate, similarly to the kernel boosting setting; see Appendix F for the formal statements. As in Theorem 4, the resulting bound exhibits a trade-off in $T$ between reducing the training smooth CE and controlling the complexity term, which grows at the rate $\mathcal{O}(\sqrt{wT})$. Similar to Corollary 4, we show that under a suitable choice of hyperparameters, both $\epsilon$-smooth CE and $\epsilon$-classification error can be guaranteed.
@@
Suppose the same assumptions as those for Eq. (26) hold. If for any $\epsilon>0$, the hyperparameters satisfy one of the following:
@@
(ii) $\beta=0$, $m=\Theta\left(\gamma^{-2}\epsilon^{-3/2}\log(1/\epsilon)\right)$, $T=\Theta\left(\gamma^{-2}\epsilon^{-1}\log^{2}(1/\epsilon)\right)$, $w=\Theta(m^{-1})$, $n=\tilde{\Omega}(\epsilon^{-2})$, then with probability at least $1-\delta$, gradient descent with the stepsize $w$ finds a parameter $\theta^{(t)}$ satisfying $\mathrm{smCE}(g_{\theta^{(t)}},S_{\mathrm{tr}})\leq\epsilon$ and $\mathbb{P}_{(X,Y)\sim\mathcal{D}}\left[(2Y-1)g_{\theta^{(t)}}(X)\leq 0\right]\leq\epsilon$ within $T$ iterations.
See Appendix F for the formal statement with explicit constants. This statement shows that both $\epsilon$-smooth CE and $\epsilon$-classification error can be achieved with $\mathcal{O}(1/\epsilon^{2})$ or $\mathcal{O}(1/\epsilon)$ iterations, assuming that the NTK-transformed data is separable with margin $\gamma$. Among the two hyperparameter settings, (ii) achieves the same error target with fewer iterations $T$ by increasing the number of hidden units $m$, and furthermore offers improved complexity compared with Corollary 4 for kernel boosting. See Appendix H for empirical results illustrating how the smooth CE and accuracy behave with $T$ and $n$.
@@ 5 Related work
To quantify the deviation from perfect calibration, various calibration metrics, including ECE, have been proposed [guo2017calibration, rahaman2021uncertainty, minderer2021revisiting]. To estimate the conditional expectation $\mathbb{E}[Y|f(X)]$ in the ECE, the binning ECE has been widely used, which partitions the probability interval into discrete bins. Despite its popularity, binning ECE suffers from being sensitive to the choices of hyperparameters, such as the number of bins, and lacks consistency and smoothness as a distance metric [kumar2019verified, nixon2019measuring, minderer2021revisiting]. A comprehensive comparison of ECE variants beyond binning is discussed by gruber2022better. blasiok2023unify systematically studied calibration distances and formalized the true distance from calibration as a rigorous notion of deviation from perfect calibration. As an efficiently estimable surrogate, they proposed the smooth CE, see also the study by foster2018smooth. A subsequent work has shown that the smooth CE can be efficiently evaluated in practice [hu2024testing]. As demonstrated by blasiok2023unify, the smooth CE also provides both upper and lower bounds on the binning ECE, making it a theoretically well-founded proxy. Therefore, the results established in this paper for the smooth CE can also be used to derive implications for the binning ECE; see Appendix B for details.
@@
Previous works have also linked the post-processing gap to the population smooth CE [blasiok2023unify, blasiok2023when]; thus, it is not readily applicable to algorithms based on the finite training dataset. Similarly, although gruber2022better discussed the post-processing, they face the same limitation. In contrast, our approach directly analyzes the smooth CE from finite data. From a generalization perspective, NEURIPS2024_futami developed algorithm-dependent bounds for binning ECE using information-theoretic techniques. In contrast, we here derive a uniform convergence bound for smooth CE, which applies more broadly beyond binning. Moreover, their focus is solely on the generalization gap, without providing insights into when the training ECE becomes small.
Although in recent studies, new boosting algorithms have been proposed to improve various notions of calibration [johnson18a, pmlr-v202-globus-harris23a], our analysis provides a theoretical explanation for why existing gradient boosting already yields strong calibration performance. Unlike prior works on boosting, which primarily focus on accuracy on the basis of functional gradients [Zhang2005BoostingWE, nitanda18a, nitanda2019gradient], we leverage functional gradients to analyze calibration, offering a novel perspective on the behavior of these algorithms.

### v4

@@ 4.1 Case study I: Gradient Boosting Tree
Under the same assumptions as Theorem 3, there exist universal constants $\{C_{i}\}$s such that, with probability at least $1-\delta$ over the draw of $S_{\mathrm{tr}}$, we have
$\mathrm{smCE}(\sigma(\bar{g}^{(T)}),\mathcal{D})\leq\frac{L_{n}(g^{(0)})}{\gamma BwT}+\frac{wB}{8\gamma}+\frac{C_{2}}{\sqrt{n}}+C_{3}wT\sqrt{\frac{2^{m}\log nd}{n}}+2\sqrt{\frac{\log(2/\delta)}{n}}.$ (19)
We find that increasing the number of steps $T$ reduces the training smooth CE, but also enlarges the function class, as the Rademacher complexity grows at rate $\mathcal{O}(wT)$. This highlights a trade-off between lowering training smooth CE and controlling model complexity. Corollary 3 shows that with appropriate hyperparameters, one can attain any target precision $\epsilon$ for the smooth CE.
In addition, bounding the norm of the functional gradient provides a generalization guarantee for test accuracy [nitanda18a]. For completeness, we include in Appendix E a formal upper bound on the misclassification rate, $P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]$.
Combining this with Theorem 2, we conclude that with suitable choices of $T$ and $n$, both the smooth CE and the misclassification rate can be controlled to within any target precision $\epsilon$.
@@
Under assumptions in Corollary 2, for any $\epsilon>0$, if the hyperparameters satisfy:
$T=\Omega(\gamma^{-2}\epsilon^{-2}),\quad w=\Theta(\gamma^{-1}\epsilon^{-1}T^{-1}),\quad n=\tilde{\Omega}(\gamma^{-2}\epsilon^{-4}),$ (20)
then, with probability at least $1-\delta$, the averaged predictor $\bar{g}^{(T)}$ satisfies $\mathrm{smCE}(\sigma(\bar{g}^{(T)}),\mathcal{D})\leq\epsilon$ and $P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq\epsilon$.
[…]
@@ 4.2 Case study II: Kernel Boosting
where $\mathcal{T}_{k}\nabla_{g}L_{n}(g)\coloneqq\frac{1}{n}\sum_{i=1}^{n}k(X_{i},\cdot)\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X_{i})),Y_{i}))$ is the empirical kernel operator. In kernel boosting, we set $h_{t}=-\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})$ in Eq. equation 14 and use the update rule:
@@
This assumption is essentially equivalent to Assumption 2 in the GBT analysis. Setting $q$ in Assumption 2 as a unit vector yields Assumption 3. Conversely, taking convex combinations of $\left(2Y_{i}-1\right)\phi(X_{i})\geq\gamma$ in Assumption 3 recovers Assumption 2.
Under Assumption 3, the $L_{1}$ norm of the functional gradient is bounded as follows:
@@
All proofs for this section are provided in Appendix F. Combining this with Eq. equation 17, we can control the smooth CE via $\left\|\mathcal{T}_{k}\nabla_{g}L_{n}(g)\right\|_{\mathcal{H}}$. Using this relation, we obtain the following result.
@@
Suppose Assumption 3, $\sup_{x,x^{\prime}\in\mathcal{X}}k(x,x^{\prime})\leq\Lambda$, and $\|g^{(0)}\|_{\mathcal{H}}\leq\Lambda^{\prime}$ hold. When using the constant stepsize $w_{t}=w$ which satisfies $w<4/\Lambda$, the average of function $\bar{g}^{(T)}=\frac{1}{T}\sum_{t=0}^{T-1}g^{(t)}$ satisfies
@@
$\mathrm{smCE}(\sigma(\bar{g}^{(T)}),\mathcal{D})\leq\frac{1}{\gamma}\sqrt{\frac{L_{n}(g^{(0)})}{wT}}\!+\frac{C_{2}}{\sqrt{n}}\!+C_{4}(\Lambda^{\prime}+\sqrt{2wTL_{n}(g^{(0)})})\sqrt{\frac{\Lambda}{n}}\!+2\sqrt{\frac{\log\frac{2}{\delta}}{n}}.$ (25)
Although covering number bounds require assumptions on kernel eigenvalue decay, here we use the Rademacher complexity bound from Theorem 2, which yields a simpler result.
We observe that increasing the number of steps $T$ decreases the training smooth CE but simultaneously enlarges the function class, since the Rademacher complexity of predictors (the third term on the right-hand side) grows with the norm of $\bar{g}^{(t)}$ at rate $\mathcal{O}(\sqrt{wT})$. This highlights a trade-off between reducing training smooth CE and controlling model complexity.
Since the norm of the functional gradient also upper bounds the misclassification rate, we show—analogous to Corollary 3—that suitable hyperparameter choices guarantee any target precision $\epsilon$ for both the smooth CE and the misclassification rate.
@@
then, with probability at least $1-\delta$, the average of function $\bar{g}^{(T)}$ satisfies $\mathrm{smCE}(\sigma(\bar{g}^{(T)}),\mathcal{D})\leq\epsilon$ and $P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq\epsilon$.
@@ 4.3 Case study III: Two-layer neural network
The next example is the two-layer neural network (NN). While NNs are typically trained by gradient descent on their parameters, recent work [nitanda2019gradient] shows that under certain hyperparameter settings—such as output scaling and a sufficiently large number of neurons—two-layer NNs behave similarly to kernel boosting with the neural tangent kernel (NTK).
We adopt the setting of nitanda2019gradient. Define the logit function $g_{\theta}:\mathcal{X}\rightarrow\mathbb{R}$ by a two-layer NN $g_{\theta}(x)=\frac{1}{m^{\beta}}\sum_{r=1}^{m}a_{r}\phi(\theta_{r}\cdot x)$, where $m>0$ is the number of hidden units, $\beta\in[0,1]$ is a scaling exponent, and $\phi:\mathbb{R}\to\mathbb{R}$ is a smooth activation function (e.g., sigmoid, tanh). The weights $\{a_{r}\}_{r=1}^{m}\in\{-1,1\}^{m}$ are fixed, while the parameters $\theta=\{\theta_{r}\}_{r=1}^{m}$ with $\theta_{r}\in\mathbb{R}^{d}$ are updated via gradient descent: $\theta^{(t+1)}=\theta^{(t)}-w\nabla_{\theta}L_{n}(g_{\theta^{(t)}})$ with constant stepsize $w>0$.
We now give an informal statement of Theorem 2 from nitanda2019gradient, which provides guarantees for the functional gradient. Assume: (1) the activation function is smooth; (2) the initial parameters are sampled from a sub-Gaussian distribution; (3) the NTK-transformed data are separable with margin $\gamma$; (4) the stepsize is sufficiently small; and (5) the number of hidden units $m$ is large enough. Then, for $T\leq\frac{m\gamma^{2}K_{3}}{w}$, there exist constants $K_{1}$, $K_{2}$, and $K_{3}$ such that, for all $\beta\in[0,1]$, with probability at least $1-\delta$ over the random initialization, the following bound holds:
@@
See Appendix G for the formal statement. Then, similar to Theorem 4, we have
@@
Combining the Rademacher complexity estimate of nitanda2019gradient with Theorem 2, we derive upper bounds on the population smooth CE and misclassification rate, analogous to the kernel boosting setting (see Appendix G for details). As in Theorem 4, the bound reveals a trade-off in $T$: increasing $T$ reduces the training smooth CE but enlarges the complexity term, which grows at rate $\mathcal{O}(\sqrt{wT})$. Similar to Corollary 4, we further show that with suitable hyperparameters, both $\epsilon$-smooth CE and $\epsilon$-classification error can be guaranteed.
@@
Suppose the same assumptions as those for Eq. equation 27 hold. If for any $\epsilon>0$, the hyperparameters satisfy one of the following:
@@
(ii) $\beta=0$, $m=\Theta\left(\gamma^{-2}\epsilon^{-3/2}\log(1/\epsilon)\right)$, $T=\Theta\left(\gamma^{-2}\epsilon^{-1}\log^{2}(1/\epsilon)\right)$, $w=\Theta(m^{-1})$, $n=\tilde{\Omega}(\epsilon^{-2})$, then with probability at least $1-\delta$, gradient descent with the stepsize $w$ finds a parameter $\theta^{(t)}$ satisfying $\mathrm{smCE}(\sigma(g_{\theta^{(t)}}),S_{\mathrm{tr}})\leq\epsilon$ and $\mathbb{P}_{(X,Y)\sim\mathcal{D}}\left[(2Y-1)g_{\theta^{(t)}}(X)\leq 0\right]\leq\epsilon$ within $T$ iterations.
See Appendix G for the formal statement with explicit constants. This result shows that both $\epsilon$-smooth CE and $\epsilon$-classification error can be achieved within $\mathcal{O}(1/\epsilon^{2})$ or $\mathcal{O}(1/\epsilon)$ iterations, assuming the NTK-transformed data are separable with margin $\gamma$. Between the two hyperparameter settings, (ii) attains the same error target with fewer iterations $T$ by increasing the number of hidden units $m$, and further yields improved complexity compared with Corollary 4 for kernel boosting. Appendix I provides empirical results illustrating how smooth CE and accuracy vary with $T$ and $n$.
@@ 5 Related work
Various metrics have been proposed to quantify deviations from perfect calibration, including ECE [guo2017calibration, rahaman2021uncertainty, minderer2021revisiting]. The binning ECE has been widely used to estimate the conditional expectation $\mathbb{E}[Y|f(X)]$ in the ECE, which partitions the probability interval into discrete bins. Despite its popularity, binning ECE is sensitive to hyperparameters such as the number of bins, and it lacks both consistency and smoothness as a distance metric [kumar2019verified, nixon2019measuring, minderer2021revisiting]. A broader comparison of ECE variants is given by gruber2022better. blasiok2023unify systematically studied calibration distances and formalized the true distance from calibration as a rigorous notion of deviation from perfect calibration. As an efficiently estimable surrogate, they proposed the smooth CE, building on foster2018smooth, kakade2008deterministic. More recently, hu2024testing showed that smooth CE can be efficiently evaluated in practice. Importantly, blasiok2023unify proved that smooth CE provides both upper and lower bounds on binning ECE, making it a theoretically sound proxy. Hence, the results established here for smooth CE also yield implications for binning ECE; see Appendix B for details.
@@
Prior work has connected the post-processing gap to the population smooth CE [blasiok2023unify, blasiok2023when], making it less applicable to algorithms trained on finite datasets. Similarly, although gruber2022better discussed post-processing, their approach faces the same limitation. In contrast, our analysis directly targets the smooth CE from finite data. From a generalization perspective, NEURIPS2024_futami developed algorithm-dependent bounds for binning ECE using information-theoretic techniques. By contrast, we derive a uniform convergence bound for smooth CE, which applies more broadly beyond binning. Moreover, their generalization bound for ECE converges at a rate $\mathcal{O}(\log n/n^{1/3})$, which is slower than our rate. We conjecture that this slower convergence arises from the nonparametric estimation of conditional probabilities via binning, whereas smooth CE avoids such estimation and hyperparameter dependence. Finally, their results focus only on the generalization gap and do not address when the training ECE itself becomes small.
Although in recent studies, new boosting algorithms have been proposed to improve various notions of calibration [johnson18a, pmlr-v202-globus-harris23a], our analysis provides a theoretical explanation for why existing gradient boosting already yields strong calibration performance. Unlike prior works on boosting, which primarily focus on accuracy on the basis of functional gradients [Zhang2005BoostingWE, nitanda18a, nitanda2019gradient], we leverage functional gradients to analyze calibration, offering a novel perspective on the behavior of these algorithms in calibration. While our analysis assumes a constant stepsize, the choice of stepsize in boosting is crucial for achieving better performance [telgarsky2013margins]. Extending our calibration analysis beyond the constant stepsize setting is an important direction for future work.